A survey found that 40​% of 822 randomly sampled teens said that their parents checked to see what Web sites they visited. The following year the same question posed to 809 teens found 42​% reporting such checks. Do these results provide evidence that more parents are​ checking?

Let p1 and p2 be the proportions of parents checking the Web sites visited in the first and second years respectively. Choose the correct null and alternative hypotheses below.

Determine the test statistic

State the result of this hypothesis. Assume alphaαequals=0.05.Choose the correct answer below
Answer #1

H0 : p1 - p2 = 0
H1 ; p1 - p2 < 0

test statistics;

p = (p1 * n1 + p2 * n2) / (n1 + n2)
= ( 0.40 * 822 + 0.42 * 0.809)/(822 + 809)
= 0.2018

SE = sqrt{ p * ( 1 - p ) * [ (1/n1) + (1/n2) ] }

= sqrt(0.2018 *(1-0.2018) *((1/822) + (1/809)))
= 0.0199

z = (p1 - p2)/SE
= ( 0.40 -0.42)/0.0199
= -1.01

p value = 0.1562

Fail to reject H0 as p value > 0.05